OpenStudy (anonymous):

Use the laws of exponents to simplify. 6^(3/4)*6^(5/8)

jimthompson5910 (jim_thompson5910):

hint: when multiplying exponential expressions with the same base, you add the exponents to get a simplified answer.

jhonyy9 (jhonyy9):

make for exponent 3/4 + 5/8

OpenStudy (anonymous):

Think about a simple example... \[X ^{2} =X*X\] \[X ^{3}=X*X*X\] \[X ^{2}*X ^{3} = (X*X)(X*X*X)\] \[X ^{2}*X ^{3} = X*X*X*X*X\] \[X ^{2}*X ^{3} = X^{5}\] All you have to do is add the exponents together. \[X^{m}*X^{n}= X^{m+n}\] get the exponents to have the same denominator. Turn 3/4 into 6/8. \[6^{3/4}*6^{5/8}= 6^{6/8}*6^{5/8}\] Then add the exponents together. \[6^{6/8}*6^{5/8}=6^{11/8}\]